OpenStudy (anonymous):

ALGEBRA 2 HELPPPP Solve the following system of equations. x + 3y + 2z = -16 2x - y + 2z = -15 2x - 2y - z = 1 (-3, -1, -5) (-3, 5, -2) (2, -4, -3) (4, 4, -1)

OpenStudy (igreen):

Okay, just plug in the solutions into each equation and see if they're true.

OpenStudy (igreen):

\(x + 3y + 2z = -16\) \( (-3, -1, -5)\) \(-3 + 3(-1) + 2(-5) = -16\) \(-3 - 3 - 10 = -16\) \(-6 - 10 = -16\) \(-16 = -16\) This is true. Now let's move onto the next equation: \(2x - y + 2z = -15\) \((-3, -1, -5)\) \(2(-3) + 1 + 2(-5) = -15\) \(-6 + 1 + 2(-5) = -15\) \(-6 + 1 - 10 = -15\) \(-5 - 10 = -15\) \(-15 = -15\) This is also true, now let's do the last equation: \(2x - 2y - z = 1\) \( (-3, -1, -5)\) \(2(-3) - 2(-1) + 5 = 1\) \(-6 - 2(-1) + 5 = 1\) \(-6 + 2 + 5 = 1\) \(-4 + 5 = 1\) \(1 = 1\) This is also true. So Option A is your answer.

OpenStudy (anonymous):

@iGreen thank you!